OpenStudy (anonymous):

Solve: 3x(x + 4) + 3(x + 4) = 0 A) 0 and -4 B) -1 and 4 C) -1 and -4 D) 1 and 4

OpenStudy (raden):

u can use distibutive property : 3x(x + 4) + 3(x + 4) = 0 (3x+3)(x+4)=0 solve for x

OpenStudy (anonymous):

B rite

hero (hero):

I don't agree with B as the answer.

OpenStudy (anonymous):

umm ok..when i put in the answers 3 times -1+3 times -4+4=0 that would make my answer cc

OpenStudy (zehanz):

You could see right away that -4 is a solution, because both terms are 0 with x = -4. Now if x is not equal to -4, you could divide everything by x-4, yielding: 3x+3=0, so 3x = -3 and x = -1. Solutions: -1 and -4. (C)

OpenStudy (anonymous):

thankzz @ZeHanz